A bag contains 3 red and 7 black balls, 2 balls are taken out at random, without replacement. If the first ball taken out is red, then what is the probability that the second taken out balll is also red?

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
Total no. of balls = 10
The probability for taking out 1 red ball = 3/10
Probability for taking out 1 red ball out of remaining ones = 2/9 .

Hence, probability for both balls to be red = 3/10 X 2/9 = 1/15

Thanks & Regards
Bharat Bajaj
askIITians Faculty
IIT Delhi